Analysis of the binary detectors

data in the 96 beam-test

Tim Dubbs (SCIPP-UCSC) Chris Hackett (UCI) Marjorie Shapiro (LBL) José Julio Lozano Bahilo (LBL)

Table of Contents

Binary detectors under test 1 Analysis scheme 2 Information on data 3 Threshold scans for different voltages 4 MPH and efficiency as function of bias voltage 10 Noise occupancy and efficiency vs. noise occupancy 12 Threshold scans varying rotation angle 15 MPH and efficiency vs. angle 18 Lorentz angle from cluster width 21 Conclusions 22

[figure]

Binary detectors under test

The binary modules from LBL installed at H8 were ATT7 and ATT8

ATT8

• Hammamatsu detector

• n^{+} strips over n bulk

• 75μm. pitch

• AC coupling

• 2 CAFEs + 2 CDPs : 256 readout channels

ATT7

• Hammamatsu detector

• n^{+} strips over n bulk

• 75μm. pitch

• AC coupling

• 2 CAFEs + 2 CDPs : 256 readout channels

Same specifications but detector and electronics irradiated separately to a fluence equivalent to 10^{14}mips/cm^{2}

In addition we used LBIC2, LBIC1 (2nd period), which are fully populated modules and UCSC2 (3 chips) as anchor planes

[figure]

Analysis scheme

The program used is an extension of the on-line software ( G. Moorhead + W. Kroeger )

Makes use of binary anchor planes instead of telescope planes ( foreseen to be used too )

1. decodification of binary information associated to channels

2. assignment of channel to strip number

3. masking of bad and hot channels

4. clusterization of hits ( >= 1 hit )

5. correlation plots produced

6. alignment file modified to match clusters for all modules

7. a) 1st period : require match inside ±4 strips between LBIC2 0 and LBIC2 1

b) 2nd period : no match required due to dead chip in LBIC2 1

8. a) 1st period : look for match with UCSC2 inside ±2 strips b) 2nd period : require match between LBIC2 0 or LBIC2 1 and UCSC2 accordig to data availability

9. build track with cluster positions and z positions of planes

10. look for cluster inside ±2 strips of extrapolated track position in ATT7 or ATT8 if no bad or hot channels inside

[figure]

Information on data

\underline{1^{st} PERIOD} Sets of runs at the following thresholds : 0.6fC, 0.8fC, 1.0fC, 1.2fC, 1.4fC, 1.7fC, 2.0fC, 2.5fC 3.0fC, 3.5fC, 4.0fC, 4.5fC for different bias voltages of ATT7 and ATT8 :

\begin{matrix}V_{AT T 7}(V) & 50 & 100 & 150 & 200 & 250 & 275 \\ V_{AT T 8}(V) & 50 & 75 & 100 & 125 & 150 & 175 \end{matrix}

to study fully and partially depleted operation of both modules \underline{2^{nd} PERIOD} Sets of runs at the following thresholds : 0.6fC, 0.8fC, 1.0fC, 1.2fC, 1.4fC, 2.0fC, 2.5fC 3.0fC, 3.5fC, 4.0fC with V_{AT T 8} = 125V and V_{AT T 7} = 250V ( usually )

Magnet was on for most of the runs for different rotation angles : −12^{0}, −8^{0}, −3^{0}, 0^{0}, 4.5^{0}, 12^{0}, 19.5^{0}, 27^{0} ( ≈ ±15^{0} of expected Lorentz angle )

[figure]

Threshold scans for different voltages

Only some relevant sets of runs have been selected During the first part of data taking some problems make data unusable V_{AT T 7} = 150V and V_{AT T 8} = 100V • low threshold points show evidence of possible oscillations

• drop of efficiency is only seen for 2 clusters of channels

• still η close to 1 at 0.8 fC for partially depleted ATT7 V_{AT T 7} = 250V and V_{AT T 8} = 150V • possible tracking problem on chip 0 (lack of statistics to determine it)

• plateau at η = 1 for some threshold values V_{AT T 7} = 275 and V_{AT T 8} = 175V • slightly improved plateau at η = 1 for ATT7

• ATT8 shows same curves

[figure]

MPH and efficiency as function of bias voltage

• ATT8 also looses a bit of efficiency below V_{AT T 8} = 110V

• good efficiency for some low threshold values in ATT7; also at intermediate bias voltages

• η > 97% for values as high as 1.4fC for ATT8 MPH : mean pulse height; η = 50% (averaged over all the channels of a chip) value of the threshold

ATT7 is close to be completely depleted in terms of response to input charges

MPH a bit low at the peak, but the MPH is averaged for all the tracks traversing also intrstrip positions

Noise occupancy and efficiency vs. noise occupancy

ATT8 fully depleted

• good behaviour showing a straight line

• noise occupancy close to 10^{−4} for 1fC

[figure]

ATT7 fully depleted ( V_{bias} = 250V )

• some oscillating noisy (not always on) channels still present

• noise iccupancy also close or lower than 10^{−4} for low thresholds ATT7 partially depleted ( V_{bias} = 150V )

October 1, 1996  Analysis of the binary ...-13  José Julio Lozano Bahilo

[figure]

Noise occupancy vs. efficiency Conclusion : good efficiency at occupancies of the order of 10^{−4} or lower for both detectors

October 1, 1996  Analysis of the binary ...-14  José Julio Lozano Bahilo

[figure]

Threshold scans varying rotation angle

Magnet was on most of the data taking time First set of runs with same conditions and at the end of first period

• angle : 0^{0}

• magnet off

• both chips show equally good response for both detectors

October 1, 1996  Analysis of the binary ...-15  José Julio Lozano Bahilo

[figure]

Angle : −3^{0} At 15^{0} outside the expected Lorent angle, efficiencies are very close to 1 in spite of low MPH

October 1, 1996  Analysis of the binary ...-16  José Julio Lozano Bahilo

[figure]

Angle : 27^{0} No data on ATT7 chip 0 and ATT8 chip 1, because of the rotation. Same comments as before apply

October 1, 1996  Analysis of the binary ...-17  José Julio Lozano Bahilo

[figure]

MPH and efficiency vs. angle

Efficiency close to 1 except for −12^{0} for a wide range of angles

October 1, 1996  Analysis of the binary ...-18  José Julio Lozano Bahilo

[figure]

In spite of the diminishing MPH, the efficiency remains very high Lorentz angle : ATT7 (0) → 10.79^{0}; ATT7 (1) → 10.38^{0} ATT8 (0) → 15.89^{0}; ATT8 (1) → 13.63^{0} Consistent values for both chips of each module are obtained; the values are smaller for the irradiated module ATT7

Lorentz angle from cluster width

The results of the fits are : ATT7 chip 0 ⇒ 8.20 ATT7 chip 1 ⇒ 10.27 ATT8 chip 0 ⇒ 11.19 ATT8 chip 1 ⇒ 10.93

Irraditated module tends to give a lower value for the Lorentz angle

[figure]

Conclusions

• Certain number of dead and noisy channels present (≈ 15 for ATT8 and ≈ 30 for ATT7)

• Operation voltages have been determined (V_{AT T 7} might be increased slightly

• Calibration not completely understood

• Very good response and noise occupancy for ATT8

• Good behaviour of irradiated ATT7 showing good capabilities